OpenStudy (anonymous):

Let u=3ai+j-4ak and v=ai+4j+2k. a) Find the value of a if u and v are orthogonal. b) if a=2 then find a vector in the opposite direction of v and four times the magnitude of v.

OpenStudy (ajprincess):

if u and v are orthogonal u.v=0

OpenStudy (ajprincess):

v=ai+4j+2k if a=2 then v=2i+4j+2k 4v=4(2i+4j+2k) since the direction is opposite it will -4(2i+4j+2k)

OpenStudy (anonymous):

dear ajpricess..the answer given in book for a) a=2 or 2/5..any idea

OpenStudy (anonymous):

for orthogonal vectors u.v=0 hence 3a^2 +4-8a=0 or 3a^2 -6a-2a+4=0 or 3a(a-2) -2(a-2)=0 or (a-2)(3a-2)=0 or a=2,2/3

OpenStudy (ajprincess):

u.v=0 So u need to find u.v. u.v=3a*a+1*4+(-4a)*2 =3a^2+4-8a 3a^2-8a+4=0 3a^2-6a-2a+4=0 3a(a-2)-2(a-2)=0 (3a-2)(a-2)=0 3a-2=0 or a-2=0 3a=2 or a=2 a=2/3

OpenStudy (anonymous):

unit vector in the direction opposite to v is -(2i+4j+2k)/sqrt(4+16+4) here a=2 hence a vector in the opposite direction of v and four times the magnitude of v is given by -4(2i+4j+2k)/sqrt(24)=-8(i+2j+k) /(2*sqrt(6)) = -4(i+2j+k)/sqrt(6)

OpenStudy (ajprincess):

vat is the answer given for b @Nesh?

OpenStudy (anonymous):

-8/sqrt24 (i+2j+k)

OpenStudy (ajprincess):

Follow @matricked's method to find b.

OpenStudy (ajprincess):

I am really sorry for messing it up. I forgot abt the idea of unit vector.

OpenStudy (anonymous):

hi matricked and ajpricess thanks for ur help..

OpenStudy (ajprincess):

yw:)

OpenStudy (anonymous):

welcome...